TheSmartOne (thesmartone):

*Solution provided by TSO* ;p If 10 OwlBucks cost 2 dollars. And I have 32 OwlBucks in 1 of my accounts, and I have 19 OwlBucks in another account. How much is all of my OwlBucks worth?

TheSmartOne (thesmartone):

Too bad you don't get any OwlBucks for answering the question :P

OpenStudy (just_one_last_goodbye):

10 owl bucks= 2 20 owl bucks= 4 30 Owl bucks= 6

OpenStudy (just_one_last_goodbye):

32 + 19 = 51 owl bucks 30 owl bucks = $6 40 owl bucks = $8 50 owl bucks = $10

OpenStudy (just_one_last_goodbye):

I said it's around $10 and some cents...

OpenStudy (just_one_last_goodbye):

My answer was an estimation.

TheSmartOne (thesmartone):

I have to repost the solution because it got cut-off.

TheSmartOne (thesmartone):

Solution: We can solve the question using a proportion. \(\sf\Large \frac{\text{10 OwlBucks}}{\text{2 Dollars}} = \frac{\text{32 + 19 OwlBucks}}{\text{x Dollars}}\) We need to cross multiply. This is the concept of cross multiplying: \(\sf\Large \frac{a}{b} = \frac{c}{d}\) Therefore, \(\sf\Large ad = bc\) So cross multiplying this: \(\sf\Large \frac{\text{10 OwlBucks}}{\text{2 Dollars}} = \frac{\text{51 OwlBucks}}{\text{x Dollars}}\)

TheSmartOne (thesmartone):

We get: \(\sf\Large 10x = 51 \times 2\) Solve that for x. \(\sf\Large 10x = 102\) Divide both sides by 10. \(\sf\Large x = 10.2\) Therefore, I have a total of 10.20 dollars. :)

ganeshie8 (ganeshie8):

\(10\text{owl} = 2\$ \implies 1= \dfrac{2\$}{10\text{owl}}\)

ganeshie8 (ganeshie8):

\((32+19) \text{owl}\) \(=(32+19) \text{owl}\times 1\) \(=(32+19) \text{owl}\times \dfrac{2\$}{10\text{ owl}}\) \(\require{cancel}=(32+19) \cancel{\text{owl}}\times \dfrac{2\$}{10\cancel{\text{ owl}}}\) \(=(32+19) \times \dfrac{2\$}{10}\)

TheSmartOne (thesmartone):

 @iambatman ;)

TheSmartOne (thesmartone):

That's another great way to solve it! @ganeshie8 Thanks for sharing another method! (:

Directrix (directrix):

How many OwlBucks is 9409 SmartCents? @TheSmartOne

TheSmartOne (thesmartone):

0. Currently, there is no relation between smartcents and owlbucks because users could medal spam themselves to get smartcents and if they got owlbucks from that, it would be unfair.

Directrix (directrix):

> users could medal spam themselves to get smartcents and if they got owlbucks from that, it would be unfair. That sounds good but makes no "cents" to me. What is the arithmetic relationship between SmartCents and medals?

TheSmartOne (thesmartone):

1 medal gets you 1 smartcent.